# Denoising Mutual Knowledge Distillation in Bi-Directional Multiple Instance Learning

### v1

@@ 1 Introduction
Multiple Instance Learning (MIL) is a widely used framework in computational pathology, particularly for Whole Slide Image (WSI) classification [1, 2, 3, 4, 5]. Due to the extremely high resolution of the WSIs, processing them as a whole is computationally expensive. Meanwhile, obtaining pixel-level or patch-level annotations is both time-consuming and expensive, which limits the scalability of fully supervised learning approaches. In the MIL framework, each WSI is partitioned into a set of smaller image patches. These patches, regarded as instances, are aggregated into a bag that shares the corresponding slide-level label. The model is trained using only bag-level labels, while the individual patches (instances) remain unlabeled. The MIL framework alleviates the reliance on fine-grained annotations, making MIL especially suitable for WSI classification.
@@
Existing MIL approaches can be broadly categorized into two types: bag-level methods and instance-level methods. Bag-level methods aggregate the features of all instances within a bag embedding to predict the bag label, without explicitly modeling the class of individual instances. Most deep learning-based MIL belong to the prior category, including the widely used attention-based MIL (ABMIL) [6] and CLAM [7]. Instance-level methods make a prediction at the instance level and then infer the bag label based on instance scores, such as DSMIL [8]. MIL methods with integrated bag and instance branches have also been proposed to utilize the strengths of both paradigms, including approaches like TransMIL [9] and DTFD-MIL [10].
@@
Both bag-level and instance-level methods have been extensively studied, each has its own advantages and drawbacks. Bag-level methods maximize the utility of the ground truth bag labels, allowing them to excel in bag classification tasks. However, the bag embedding provably contains less information than the whole set of instance features [11]. As a direct consequence, the bag-level methods tend to focus on learning only easy instances [12] or spurious correlations [13] within a bag, limiting their generalization ability. Instance-level methods consider all instances individually. However, the lack of ground-truth supervision for individual instances makes them susceptible to noise in the instance pseudo labels during training. These drawbacks in both categories compromise the performance and robustness of MIL models.
@@
To address the issues of not exploiting sufficient instances in bag-level methods, hard instance mining (HIM) modules are proposed to mask easy instances during training [12, 14]. HIM modules encourage the model to focus on hard instances during training, thereby enhancing its ability to generalize. The IBMIL model [13] and the CAMIL model [15] are proposed recently to avoid associating bag classifiers with spurious features by eliminating harmful confounders using causal inference. The CausalMIL model [16, 17] utilizes causal inference in MIL to reduce its sensitivity to data distribution, finding stable instances to guide bag-level classification.
@@
Previous work has also investigated methods to improve the reliability of the instance-level pseudo-labels. The CIMIL model [18] used counterfactual inference and a hierarchical instance searching module to search for reliable instances and generate precise pseudolables. Optimal transport is also used to calibrate soft labels in a recent paper [11]. Qu et al. [19] use prototype learning to improve the reliability of pseudolable based on representation learned by the the so-called instance-level weakly supervised contrastive learning.
@@
Although current improvement methods have demonstrated some effectiveness, they may also introduce biases from other sources into the model. HIM may inadvertently introduce noise into the bag label when the masks filter out all positive instances, as discussed in prior work [20]. Causal inference methods depend on the accurate modeling of confounders, without which the inferred effects may be invalid [21]. In addition, some MIL frameworks adopt complex mechanisms to identify representative instances within a bag, resulting in increased implementation difficulty and reduced interpretability [22]. The use of combinatorial or multi-stage algorithms further increases computational cost, making such methods less scalable in real-world applications [23].
@@
Without external guidance such as instance level annotation, leveraging the model’s own capability to correct pseudo-label noise seems inevitable [24]. The weak to strong generalization phenomenon describes the ability of a student to outperform weak teachers in terms of generalization capability when it is pretrained to generate robust data representations [25]. Our denoising method is based on this simple idea of amplifying generalization power using only weak supervision.
@@
  • We propose a dual-level MIL architecture that include both bag branch and instance branch. Compared with previous methods, the supervision signals in both branches are enhanced: (1) the soft labels are anchored to specific classes to provide improved supervision for instance branch; (2) the instance predictions are aggregated to assist the training of the bag branch as a supplement to bag labels. Besides, the training process of the two branches are scheduled to ensure that they are optimized at the same pace.
@@ 2 Related Work
Deep multiple instance learning (Deep MIL) extends the classical MIL framework by integrating neural networks to jointly learn instance-level and bag-level representations. Wang et al. [26] proposed mi-Net and MI-Net as basic deep learning-based instance-level and embedding-level pooling methods. Ilse et al. [6] introduced the Attention-Based MIL model (ABMIL), which employs a learnable attention mechanism to aggregate instance embeddings within each bag. More recent work has used transformers [9] and graph neural networks [27, 28, 29] to aggregate instance features.
@@
Instance classifiers are incorporated in many MIL models to ensure that bag predictions are consistent with key instance predictions. An instance classifier is a component in MIL frameworks that assigns labels or scores to individual instances, enabling finer-grained modeling under weak supervision. The CLAM model [30] uses an instance classification head to ensure that the top $K$ attended patches share the same class with the bag. The DGMIL model [31] uses instance-level pseudo-labels to refine the feature distribution to separate positive and negative instance representation. Ma et al. [11] treated MIL as a semi-supervised task, training an instance classifier with pseudo labels and consistency regularization. Qu et al. [19] leveraged contrastive and prototype-based learning to enhance the discriminative ability of instance classifiers. Lin et al. [18] proposed a counterfactual inference framework named CIMIL that uses instance representation to generate more informative representations for MIL models.
@@
Hard instance mining has been widely adopted to enhance model robustness and generalization in scenarios involving ambiguous or weakly informative samples. In object detection, Shrivastava et al. [32] proposed Online Hard Example Mining (OHEM), which prioritizes training on samples with the highest loss to accelerate convergence and improve detection accuracy. This approach laid the groundwork for subsequent applications of hard mining in other domains.
@@
In the context of applying MIL for WSI classification, the WENO model [12] uses a hard positive instance mining (HPM) module to mask the most attended patches in bag classification training to force the model to learn hard instances. Tang et al. [14] introduced a masked hard instance mining (MHIM) mechanism that identifies challenging instances by enforcing instance-level consistency within a teacher-student framework. Their methods improves MIL performance by explicitly guiding the model to focus on informative but difficult regions in histopathology images, outperforming previous attention-based MIL approaches. Other work leveraging HIM includes HPA-MIL [33] and ACMIL [34].
@@
Weak-to-Strong (W2S) generalization studies methods to supervise a stronger model using supervision from a weaker model. Burns et al. [35] introduced this setting and demonstrated it empirically through several tasks. Charikar et al. [36] provided a theoretical framework quantifying how strong models can benefit from weak labels. Lang et al. [37] proposed mechanisms explaining how label correction and coverage expansion support generalization beyond the supervisor.
@@
From the application perspective, Guo et al. [38] showed that vision foundation models trained with weak supervision can outperform both their weak supervisors and fully supervised baselines. Wang et al. [39] utilize confident subset of pseudo-labels to guide the instruction tuning of LLMs. Cui et al. [40] used aggregated weak feedback to fine-tune large language models on text classification and generation tasks, achieving higher-quality outputs. Somerstep et al. [41] proposed a refinement-based transfer learning approach that enables stronger models to generalize beyond weak teachers in alignment tasks.
@@ 3.2 MIL in WSI Classification
Formally, given a bag $X_{i}=\{x_{i}^{j}\}_{j=1}^{n_{i}}$, each instance $x_{i}^{j}$ is encoded into a feature vector $h_{i}^{j}=\mathcal{E}(x_{i}^{j})\in\mathbb{R}^{d}$. An attention mechanism then computes a scalar attention score $a_{i,j}\in[0,1]$ for each instance $x_{i}^{j}$. In the ABMIL model, $a_{i,j}$ is defined as $a_{i}^{j}=w^{\top}\tanh{(V(h_{i}^{j})^{T})}$ ), where $V\in\mathbb{R}^{l\times d}$ and $w\in\mathbb{R}^{l}$ are learnable parameters. The vector $a_{i}=[a_{i}^{1},\dots,a_{i}^{n_{i}}]$ is then normalized by using a softmax function $\alpha_{i}^{j}=\nicefrac{{\text{exp}({a_{i}^{j}})}}{{\sum_{i=1}^{n_{i}}\text{exp}({a_{i}^{j}})}}$. The bag-level representation is obtained as a weighted average of instance features: $H_{i}=\sum_{j=1}^{n_{i}}\alpha_{i}^{j}\cdot\mathcal{E}(x_{i}^{j})$. This representation is then passed to a classifier $\varphi$ to predict the bag label.
@@ 3.6 Loss Functions
$L_{\text{inst}}=CE\left(\widehat{Y}_{i},\hat{y}_{i}^{j}\right)\,,\text{ where }\hat{y}_{i}^{j}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\max\limits_{j\in M_{i}}\left(\hat{y}_{i}^{j}\right)$ (9)
@@ 4.1 Datasets
To comprehensively evaluate the performance of our method, we conducted experiments on two real-world datasets. CAMELYON16 [42] is a benchmark dataset for evaluating metastasis detection algorithms, consisting of $398$ WSIs of lymph node sections stained with hematoxylin and eosin (H&E), $269$ of which are in the training set and $129$ are in the test set. Besides slide labels, it also provides annotated contours indicating ground-truth patch labels. A patch is labeled positive if at least 25% of its area is annotated as lesion tissue. Instance labels are used exclusively for testing and remain inaccessible during training. We conduct experiments on the dataset under two settings: (1) using the original training and test sets directly for training and evaluation, and (2) creating a validation set by extracting 54 WSIs from the original training set, yielding a split of 215 WSIs for training, 54 for validation, and 129 for testing. We denote these two settings as C16 and C16v, respectively.
TCGA-NSCLC [^1] Official website for TCGA: https://www.cancer.gov/ccg/research/genome-sequencing/tcga has two cancer subtypes, lung adenocarcinoma (LUAD) and lung squamous cell carcinoma (LUSC). The dataset contains $478$ LUAD slides and 478 LUSC slides, with only slide labels provided. All experiments on TCGA-NSCLC are conducted using $4$-fold cross-validation, with the data split into training, validation, and test sets in a ratio of 13:2:5.
@@ 4.2 Evaluation Metrics and Implementation Details
Implementation Details: Each WSI slide is cropped into a series of non-overlapping patches with a size of 512 $\times$ 512, and background regions are removed. A ResNet-50 network pretrained on the ImageNet is employed to extract a 1024-dimensional embedding vector from each patch for the training step. The shared encoder is a fully connected layer followed by a non-linear module. During training, we use Adam optimizer with a learning rate of $2\times 10^{-4}$ and weight decay of $1\times 10^{-5}$. The mini-batch sizes of the bag and the instance branch are $1$ and $1024$, respectively. The bag-level training and instance-level training are scheduled to run periodically. In each cycle, the algorithm run $10$ epochs of bag-level training, followed by one epoch of instance-level training. All models are trained for $200$ epochs on CAMELYON16 in the C16 setting, and at most $200$ epochs with early stopping in the C16v setting. For TCGA-NSCLC, early stopping is applied during training, with a maximum of $200$ epochs and patience set to $20$. All experiments were performed on a machine with 8 NVIDIA GeForce RTX 3090 GPUs.
@@ 4.3 Baselines
We compare our method with existing methods, including ABMIL, DSMIL, CLAM, and TransMIL. We also implement existing end-to-end boosting frameworks, WENO and MHIM-MIL. All boosting frameworks are tested by combining with ABMIL and DSMIL, respectively. We reproduced the baseline methods based on their published code [^2] A unified framework for MIL methods: https://github.com/lingxitong/MIL_BASELINE (under the MIT license); official implementation of WENO: https://github.com/miccaiif/WENO (unlicensed); official implementation of MHIM-MIL: https://github.com/DearCaat/MHIM-MIL (unlicensed)., with necessary modifications to adapt to different dataset settings. We exclude CIMIL and CausalMIL from the baselines since they have multi-stage training processes, which significantly increases computational and implementation complexity. For methods without instance-level prediction output, we find a reasonable approach to equip the model with such a capability and report the result.
@@ 4.4 Results
Table 1 shows the performance of binary classification between positive and negative on CAMELYON16 and subtyping between LUAD and LUSC on TCGA-NSCLC. Results of all models are presented, except for the fully supervised method on TCGA-NSCLC due to the lack of instance labels, and the instance-level results of TransMIL on CAMELYON16 since the features of instances are fused in several modules and therefore no instance prediction is available.
@@
Figure 2: Visualization of instance-level prediction results. (a) shows one original CAMELYON16 slide thumbnail with ground-truth positive instances masked in green. The following are heatmaps overlapped on the slide image where deeper red indicates higher predicted probability of being positive and deeper blue indicates lower negative probability.
Ablation Study: To examine the effect of adding self-confidence loss to both the bag branch and the instance branch, we conduct ablation study on the CAMELYON16 dataset (in the C16 setting). We use the ABMIL model in a modified version of our method where $L_{\text{self}}$ and $\ell_{\text{self}}$ are removed from the loss functions. The bag-level (instance-level) AUC drops from 0.9318 (0.9216) to 0.9151 (0.9117), with a reduction of 0.0167 (0.0099). The bag-level (instance-level) ACC decreased from 0.9173 (0.9291) to 0.8915 (0.9149), with a drop of 0.0258 (0.0142).

### v2

@@ 1 Introduction
Multiple Instance Learning (MIL) is a widely used framework in computational pathology, particularly for Whole Slide Image (WSI) classification 10.1007/978-3-030-87237-3_32 ; campanella2019clinical ; SRINIDHI2021101813 ; 9975198 ; Qu_2022 . Due to the extremely high resolution of the WSIs, processing them as a whole is computationally expensive. Meanwhile, obtaining pixel-level or patch-level annotations is both time-consuming and expensive, which limits the scalability of fully supervised learning approaches. In the MIL framework, each WSI is partitioned into a set of smaller image patches. These patches, regarded as instances, are aggregated into a bag that shares the corresponding slide-level label. The model is trained using only bag-level labels, while the individual patches (instances) remain unlabeled. The MIL framework alleviates the reliance on fine-grained annotations, making MIL especially suitable for WSI classification.
@@
Existing MIL approaches can be broadly categorized into two types: bag-level methods and instance-level methods. Bag-level methods aggregate the features of all instances within a bag embedding to predict the bag label, without explicitly modeling the class of individual instances. Most deep learning-based MIL belong to the prior category, including the widely used attention-based MIL (ABMIL) ilse2018attention and CLAM clam . Instance-level methods make a prediction at the instance level and then infer the bag label based on instance scores, such as DSMIL li2021DSMIL . MIL methods with integrated bag and instance branches have also been proposed to utilize the strengths of both paradigms, including approaches like TransMIL shao2021transmil and DTFD-MIL DTFD .
@@
Both bag-level and instance-level methods have been extensively studied, each has its own advantages and drawbacks. Bag-level methods maximize the utility of the ground truth bag labels, allowing them to excel in bag classification tasks. However, the bag embedding provably contains less information than the whole set of instance features ma2024rethinkingmultipleinstancelearning . As a direct consequence, the bag-level methods tend to focus on learning only easy instances qu2022WENO or spurious correlations Lin_2023_CVPR within a bag, limiting their generalization ability. Instance-level methods consider all instances individually. However, the lack of ground-truth supervision for individual instances makes them susceptible to noise in the instance pseudo labels during training. These drawbacks in both categories compromise the performance and robustness of MIL models.
@@
To address the issues of not exploiting sufficient instances in bag-level methods, hard instance mining (HIM) modules are proposed to mask easy instances during training qu2022WENO ; tang2023multiple . HIM modules encourage the model to focus on hard instances during training, thereby enhancing its ability to generalize. The IBMIL model Lin_2023_CVPR and the CAMIL model Chen_Sun_Zhao_2024 are proposed recently to avoid associating bag classifiers with spurious features by eliminating harmful confounders using causal inference. The CausalMIL model DBLP:conf/ecai/ZhangLL20 ; NEURIPS2022_e261e92e utilizes causal inference in MIL to reduce its sensitivity to data distribution, finding stable instances to guide bag-level classification.
@@
Previous work has also investigated methods to improve the reliability of the instance-level pseudo-labels. The CIMIL model lin2024boosting used counterfactual inference and a hierarchical instance searching module to search for reliable instances and generate precise pseudolables. Optimal transport is also used to calibrate soft labels in a recent paper ma2024rethinkingmultipleinstancelearning . Qu et al. qu2024rethinking use prototype learning to improve the reliability of pseudolable based on representation learned by the the so-called instance-level weakly supervised contrastive learning.
@@
Although current improvement methods have demonstrated some effectiveness, they may also introduce biases from other sources into the model. HIM may inadvertently introduce noise into the bag label when the masks filter out all positive instances, as discussed in prior work 7298968 . Causal inference methods depend on the accurate modeling of confounders, without which the inferred effects may be invalid 9363924 . In addition, some MIL frameworks adopt complex mechanisms to identify representative instances within a bag, resulting in increased implementation difficulty and reduced interpretability 7792710 . The use of combinatorial or multi-stage algorithms further increases computational cost, making such methods less scalable in real-world applications 7780882 .
@@
Without external guidance such as instance level annotation, leveraging the model’s own capability to correct pseudo-label noise seems inevitable DBLP:journals/corr/abs-1908-02983 . The weak to strong generalization phenomenon describes the ability of a student to outperform weak teachers in terms of generalization capability when it is pretrained to generate robust data representations 9156610 . Our denoising method is based on this simple idea of amplifying generalization power using only weak supervision.
@@
  • We propose a dual-level MIL architecture that includes both bag branch and instance branch. Compared with previous methods, the supervision signals in both branches are enhanced: (1) the soft labels are anchored to specific classes to provide improved supervision for instance branch; (2) the instance predictions are aggregated to assist the training of the bag branch as a supplement to bag labels. Besides, the training process of the two branches are scheduled to ensure that they are optimized at the same pace.
@@ 2 Related Work
Deep multiple instance learning (Deep MIL) extends the classical MIL framework by integrating neural networks to jointly learn instance-level and bag-level representations. Wang et al. WANG201815 proposed mi-Net and MI-Net as basic deep learning-based instance-level and embedding-level pooling methods. Ilse et al. ilse2018attention introduced the Attention-Based MIL model (ABMIL), which employs a learnable attention mechanism to aggregate instance embeddings within each bag. More recent work has used transformers shao2021transmil and graph neural networks 10.1007/978-3-030-87237-3_33 ; 9779215 ; Li_2024_CVPR to aggregate instance features.
@@
Instance classifiers are incorporated in many MIL models to ensure that bag predictions are consistent with key instance predictions. An instance classifier is a component in MIL frameworks that assigns labels or scores to individual instances, enabling finer-grained modeling under weak supervision. The CLAM model lu2021data uses an instance classification head to ensure that the top $K$ attended patches share the same class with the bag. The DGMIL model 10.1007/978-3-031-16434-7_3 uses instance-level pseudo-labels to refine the feature distribution to separate positive and negative instance representation. Ma et al. ma2024rethinkingmultipleinstancelearning treated MIL as a semi-supervised task, training an instance classifier with pseudo labels and consistency regularization. Qu et al. qu2024rethinking leveraged contrastive and prototype-based learning to enhance the discriminative ability of instance classifiers. Lin et al. lin2024boosting proposed a counterfactual inference framework named CIMIL that uses instance representation to generate more informative representations for MIL models.
@@
Hard instance mining has been widely adopted to enhance model robustness and generalization in scenarios involving ambiguous or weakly informative samples. In object detection, Shrivastava et al. shrivastava2016training proposed Online Hard Example Mining (OHEM), which prioritizes training on samples with the highest loss to accelerate convergence and improve detection accuracy. This approach laid the groundwork for subsequent applications of hard mining in other domains.
@@
In the context of applying MIL for WSI classification, the WENO model qu2022WENO uses a hard positive instance mining (HPM) module to mask the most attended patches in bag classification training to force the model to learn hard instances. Tang et al. tang2023multiple introduced a masked hard instance mining (MHIM) mechanism that identifies challenging instances by enforcing instance-level consistency within a teacher-student framework. Their methods improves MIL performance by explicitly guiding the model to focus on informative but difficult regions in histopathology images, outperforming previous attention-based MIL approaches. Other work leveraging HIM includes HPA-MIL DGXLX+24 and ACMIL zhang2024attention .
@@
Weak-to-Strong (W2S) generalization studies methods to supervise a stronger model using supervision from a weaker model. Burns et al. burns2023weak introduced this setting and demonstrated it empirically through several tasks. Charikar et al. charikar2024quantifying provided a theoretical framework quantifying how strong models can benefit from weak labels. Lang et al. lang2024theoretical proposed mechanisms explaining how label correction and coverage expansion support generalization beyond the supervisor.
@@
From the application perspective, Guo et al. GCWH+24 showed that vision foundation models trained with weak supervision can outperform both their weak supervisors and fully supervised baselines. Wang et al. wang-etal-2023-self-instruct utilize confident subset of pseudo-labels to guide the instruction tuning of LLMs. Cui et al. cui2025bayesian used aggregated weak feedback to fine-tune large language models on text classification and generation tasks, achieving higher-quality outputs. Somerstep et al. somerstep2025a proposed a refinement-based transfer learning approach that enables stronger models to generalize beyond weak teachers in alignment tasks.
@@ 3.2 MIL in WSI Classification
Formally, given a bag $X_{i}=\{x_{i}^{j}\}_{j=1}^{n_{i}}$, each instance $x_{i}^{j}$ is encoded into a feature vector $h_{i}^{j}=\mathcal{E}(x_{i}^{j})\in\mathbb{R}^{d}$. An attention mechanism then computes a scalar attention score $a_{i,j}\in[0,1]$ for each instance $x_{i}^{j}$. In the ABMIL model, $a_{i,j}$ is defined as $a_{i}^{j}=w^{\top}\tanh{(Vh_{i}^{j})}$ ), where $V\in\mathbb{R}^{l\times d}$ and $w\in\mathbb{R}^{l}$ are learnable parameters. The vector $a_{i}=[a_{i}^{1},\dots,a_{i}^{n_{i}}]$ is then normalized by using a softmax function $\alpha_{i}^{j}=\nicefrac{{\text{exp}({a_{i}^{j}})}}{{\sum_{i=1}^{n_{i}}\text{exp}({a_{i}^{j}})}}$. The bag-level representation is obtained as a weighted average of instance features: $H_{i}=\sum_{j=1}^{n_{i}}\alpha_{i}^{j}\cdot\mathcal{E}(x_{i}^{j})$. This representation is then passed to a classifier $\varphi$ to predict the bag label. The definitions of attention scores vary in different models, explained in Appendix A.1.
@@ 3.6 Loss Functions
$L_{\text{inst}}=CE\left(\widehat{Y}_{i},\hat{y}_{i}\right)\,,\text{ where }\hat{y}_{i}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\max\limits_{j\in M_{i}}\left(\hat{y}_{i}^{j}\right)$ (9)
@@ 4.1 Datasets
To comprehensively evaluate the performance of our method, we conducted experiments on two real-world datasets. CAMELYON16 10.1001/jama.2017.14585 is a benchmark dataset for evaluating metastasis detection algorithms, consisting of $398$ WSIs of lymph node sections stained with hematoxylin and eosin (H&E), $269$ of which are in the training set and $129$ are in the test set. Besides slide labels, it also provides annotated contours indicating ground-truth patch labels. A patch is labeled positive if at least 25% of its area is annotated as lesion tissue. Instance labels are used exclusively for testing and remain inaccessible during training. We conduct experiments on the dataset under two settings: (1) using the original training and test sets directly for training and evaluation, and (2) creating a validation set by extracting 54 WSIs from the original training set, yielding a split of 215 WSIs for training, 54 for validation, and 129 for testing. We denote these two settings as C16 and C16v, respectively.
TCGA-NSCLC [^1] Official website for TCGA: https://www.cancer.gov/ccg/research/genome-sequencing/tcga. has two cancer subtypes, lung adenocarcinoma (LUAD) and lung squamous cell carcinoma (LUSC). The dataset contains $478$ LUAD slides and 478 LUSC slides, with only slide labels provided. All experiments on TCGA-NSCLC are conducted using $4$-fold cross-validation, with the data split into training, validation, and test sets in a ratio of 13:2:5.
@@ 4.2 Evaluation Metrics and Implementation Details
Implementation Details: Each WSI slide is cropped into a series of non-overlapping patches with a size of 512 $\times$ 512, and background regions are removed. A ResNet-50 network pretrained on the ImageNet is employed to extract a 1024-dimensional embedding vector from each patch for the training step. All models are trained for $200$ epochs on CAMELYON16 in the C16 setting, and at most $200$ epochs with early stopping in the C16v setting. For TCGA-NSCLC, early stopping is applied during training, with a maximum of $200$ epochs and patience set to $20$. All experiments were performed on a machine with 8 NVIDIA GeForce RTX 3090 GPUs. More implementation details are provided in Appendix A.3.
@@ 4.3 Baselines
We compare our method with existing methods, including ABMIL, DSMIL, CLAM, and TransMIL. We also implement existing end-to-end boosting frameworks, WENO and MHIM-MIL. All boosting frameworks are tested by combining with ABMIL and DSMIL, respectively. We reproduced the baseline methods based on their published code [^2] A unified framework for MIL methods: https://github.com/lingxitong/MIL_BASELINE (under the MIT license); official implementation of WENO: https://github.com/miccaiif/WENO (unlicensed); official implementation of MHIM-MIL: https://github.com/DearCaat/MHIM-MIL (unlicensed)., with necessary modifications to adapt to different dataset settings. We exclude CIMIL and CausalMIL from the baselines since they have multi-stage training processes, which significantly increases computational and implementation complexity. For methods without instance-level prediction output, we find a reasonable approach to equip the model with such a capability and report the result. Details about instance prediction implementations are shown in Appendix A.2.
@@ 4.4 Results
Table 1 shows the performance of binary classification between positive and negative on CAMELYON16 and subtyping between LUAD and LUSC on TCGA-NSCLC. Results of all models are presented, except for the fully supervised method on TCGA-NSCLC due to the lack of instance labels, and the instance-level results of TransMIL on CAMELYON16 since the features of instances are fused in several modules and therefore no instance prediction is available. More experimental results are provided in Appendix B.
@@
Figure 2: Visualization of instance-level prediction results. (a) shows one original CAMELYON16 slide thumbnail with ground-truth positive instances masked in green. The following are heatmaps overlapped on the slide image where deeper red indicates higher predicted probability of being positive and deeper blue indicates lower probability of being positive.
Ablation Study: To examine the effect of adding self-confidence loss to both the bag branch and the instance branch, we conduct ablation study on the CAMELYON16 dataset (in the C16 setting). We use the ABMIL model in a modified version of our method where $L_{\text{self}}$ and $\ell_{\text{self}}$ are removed from the loss functions. The bag-level (instance-level) AUC drops from 0.9318 (0.9216) to 0.9151 (0.9117), with a reduction of 0.0167 (0.0099). The bag-level (instance-level) ACC decreased from 0.9173 (0.9291) to 0.8915 (0.9149), with a drop of 0.0258 (0.0142). More ablation experiments are provided in Appendix C.